OpenStudy (calculusxy):

If a man runs 'M' miles in 'T' hours, his speed is A. M/T B. M + T C. M - T D. MT

OpenStudy (anonymous):

If he runs 10 miles in 10 hours then his speed would be 1 mile per hour. so 10/10 or M/T

OpenStudy (calculusxy):

Using "1 mile per hour", you can find the speed? So is there a formula that you can prove pied?

OpenStudy (calculusxy):

*provide

OpenStudy (anonymous):

I was just using that as an example, but you can use the formula speed=M/T for any numbers. 20 miles in 2 hours; apply the formula s=m/t; s=20/2; s=10mph

OpenStudy (calculusxy):

Can you help me with another one? I have the answer but I just would like to know if I am correct. :)

OpenStudy (anonymous):

ok

OpenStudy (calculusxy):

In how much less time does a runner who finishes a marathon in 2 hours 12 minutes 38 seconds complete the race than a runner who finishes in 3 hours 2 minutes 24 seconds?

OpenStudy (calculusxy):

A. 48 min. 56 sec. B. 49 min. 46 sec. C. 1 hr. 51 min. 22 sec. D. 1 hr. 26 min. 12 sec. I got B.

OpenStudy (anonymous):

So you would just subtract the first runners time from the second runners time: 3:02:24 -2:12:38 ---------- 0:49:46

OpenStudy (anonymous):

you just have to remember when you carry the minute over to the seconds that you are adding 60. Same goes for the hours the minutes.

OpenStudy (calculusxy):

CB=30 miles BA=40 miles Two drivers begin at point C simultaneously. One drives from C to B to A. The other drives directly to A at 50 mph. How fast must the first person drive to get to A first?

OpenStudy (anonymous):

Is this a triangle or a straight line?

OpenStudy (calculusxy):

Triangle

OpenStudy (anonymous):

gotcha. A right triangle I assume?

OpenStudy (calculusxy):

Yes

OpenStudy (anonymous):

 Like this?

OpenStudy (calculusxy):

Switch the places of the numbers.

OpenStudy (anonymous):

Ok. So using Pythagorean theorem a^2+b^2=c^2 we know that the hypotenuse is 50 miles ((900+1600=2500) sqrt of 2500 is 50) so if driver 2 is driving 50 miles at 50 miles per hour it will take him one hour to get there. Driver 1 has to drive 30 miles + 40 miles in 1 hour, so he has to drive 70 miles per hour to get there at the same time.

OpenStudy (anonymous):

so to get there first he must drive >70mph

OpenStudy (calculusxy):

How did you get more than 70 mph?

OpenStudy (anonymous):

well he has to cover the 30 miles from c to b, and then 40 miles from b to a. So he has to drive 70 miles. And if he only has 1 hour to drive 70 miles he has to drive faster than 70mph to get there first

OpenStudy (calculusxy):

Your a math wiz!

OpenStudy (anonymous):

Lol. I'm working on my masters to be a math teacher

OpenStudy (calculusxy):

Event A occurs every 4 years, event B every 11 years, and event C every 33 years. IF they last occurred together in 1950, when will they next occur simultaneously?

OpenStudy (calculusxy):

(F) 3,402 (G) 1,983 (H) 2,082 (J) 6,804

OpenStudy (calculusxy):

@quentinloy

OpenStudy (anonymous):

Sorry, what's the question?

OpenStudy (anonymous):

oh, I see it

OpenStudy (anonymous):

so 11 is a multiple of 33, so you really need to know when it will when event A and event C will coincide, but cause whenever C occurs, so does B. So event C occurs in 33, 66, 99, 132 years. A does not occur in 33, 66, or 99, but it does occur in 132. 4x33=132. 1950+132 years is when all three events will occur at the same time